OpenStudy (liv1234):

A rare bacterial culture is being grown in a lab. As the days progress, the cells multiply and grow. After 1 day, there is only 1 cell. After 2 days, there are 8 cells. After 3 days, there are 18 cells. After 4 days, there 31 cells. Which recursive equation represents the pattern? A. an=an-1+3n+1 <- my answer B. an=an-1+3(n+1) C. an=an-1+(n+3) D. an=an-1+(n-3)

OpenStudy (lfghjyt):

Substituting the values for day 4, 18 + 3*4 + 1 = 31 18 + 3(4 + 1)= 33 18 + (4 + 3) = 25 18 + (4 - 3) = 19 Thus the equation is, a˅n = a˅(n–1) + 3*n + 1 .... ˅ means sub

OpenStudy (liv1234):

@lfghjyt Does that mean I'm right?

OpenStudy (liv1234):

@sissyedgar

OpenStudy (liv1234):

@inkyvoyd

OpenStudy (sissyedgar):

I think it does but i'm not sure. It looks like it to me tho.